# Teaching Transformers to Solve Combinatorial Problems through Efficient Trial & Error

### v1

@@ Abstract
Despite their proficiency in various language tasks, Large Language Models (LLMs) struggle with combinatorial problems like Satisfiability, Traveling Salesman Problem, or even basic arithmetic. We address this gap through a novel approach for solving problems in the class NP. We focus on the paradigmatic task of Sudoku and achieve state-of-the-art accuracy (99%) compared to prior neuro-symbolic approaches. Unlike prior work that used custom architectures, our method employs a vanilla decoder-only Transformer (GPT-2) without external tools or function calling. Our method integrates imitation learning of simple Sudoku rules with an explicit Depth-First Search (DFS) exploration strategy involving informed guessing and backtracking. Moving beyond imitation learning, we seek to minimize the number of guesses until reaching a solution. We provide a rigorous analysis of this setup formalizing its connection to a contextual variant of Min–Sum Set Cover, a well-studied problem in algorithms and stochastic optimization.
@@ 1 Introduction
Large Language Models (LLMs) based on the Transformer architecture have shown remarkable performance in a wide range of Machine Learning tasks, achieving state-of-the-art results on language understanding and generation benchmarks (VSP+, 17; BMR+, 20; RWC+, 19). Trained to predict the next token in large text corpora, these models acquire strong reasoning abilities, but whether they perform genuine logical reasoning or rely on pattern-matching remains under debate (VMO+, 23; DLS+, 23). Prompting strategies such as Chain-of-Thought and its many variants (NAGA+, 21; WWS+, 22; YYZ+, 23; WWS+, 23; BBK+, 24; CMWC, 23; KTF+, 22; ZSH+, 23; MTG+, 23) have enabled LLMs to exhibit multi-step reasoning capabilities and solve even complex mathematical problems (HBK+, 21). Yet, LLMs still struggle in more structured domains like arithmetic and combinatorial optimization. Controlled evaluations on synthetic tasks, including algorithmic puzzles, formal grammars, logic-based games, and arithmetic problems, have revealed consistent failure modes of current models and methods (LAG+, 22; LHB+, 23; NLW, 23; AZL, 23; YXLAZ24a, ; YXLAZ24b, ; LSL+, 24). At the same time, the literature highlights that Transformer-based models have the potential to leverage internal mechanisms like attention to capture combinatorial structure without specialized heuristics or external tools. However, appropriately training them to display this behavior remains a major challenge and an active topic of research in robustness and generalization.
@@
Our primary testbed in this work is Sudoku, chosen for its well-defined rules, extensively studied difficulty levels, and widespread use as a reasoning benchmark. A standard Sudoku puzzle consists of a $9\times 9$ grid, subdivided into nine $3\times 3$ boxes. The grid is partially filled with digits from 1 to 9, and the goal is to complete it subject to the following rules: (i) each cell contains exactly one digit, (ii) each row contains every digit exactly once, (iii) each column contains every digit exactly once, and (iv) each box contains every digit exactly once. We focus on the standard Sudoku puzzles, while our approach can be directly applied to its generalized form with $n\times n$ grids which is NP-complete (YS, 03) and thus equivalent to any problem in the class NP.
Solving Sudoku puzzles without external tools is extremely challenging even for state-of-the-art industrial-scale LLMs. We evaluated several models, including OpenAI GPT (AAA+, 23), Google Gemini (TAB+, 23), and DeepSeek (GYZ+, 25), on randomly generated Sudoku puzzles. As shown in Table 1, none of these models were able to produce correct solutions. We tested both general-purpose LLMs and reasoning-focused variants, since Sudoku demands logical reasoning. General-purpose models generated answers quickly, whereas reasoning-oriented models ran significantly longer as they attempted step-by-step strategies mirroring human solving patterns. However, once they made an incorrect inference, these models could not backtrack and revise their solutions, often ending with contradictory board states. Additional details are provided in Appendix B.
@@
Beyond these general-purpose models, several methods have been proposed in the literature to solve Sudoku puzzles via Deep Learning although often relying on specialized architectures. (MKPZ, 11) investigates a Neural Network (NN)-based approach based on Hopfield Networks as an integer optimization tool, addressing its convergence limitations through a coprocessor NN designed to solve linear programming problems. (PPW, 18) introduces the Recurrent Relational Network (RRN), a generalization of the Relational Network (SRB+, 17), to evaluate multi-step relational reasoning. Their method involves multiple rounds of message passing between Sudoku puzzle cells to arrive at a solution. (YIL, 23) explores solving Sudoku puzzles using a Recurrent Transformer model, integrating Sudoku’s constraints directly into both the model architecture and the training process.
@@
Closer to our work, (SDWP, 24) study how a causal language model (GPT-based) can reliably solve Sudoku puzzles. By reframing Sudoku as a next-token prediction task over transcripts of human-like reasoning based on known Sudoku strategies, their Causal Transformer learns to imitate and apply these heuristics incrementally, typically executing simpler strategies before tackling more complex ones. This approach showcases the promising capability of Autoregressive Transformers to successfully internalize and deploy multi-step reasoning procedures in a more interpretable way. However, their proposed method trains the Transformer on a structured sequence of logical steps, which do not allow the model to make mistakes. This inherently limits its problem-solving capability to strategies present in the training data whereas a large fraction of puzzles are not solvable by simple logic steps. As such, its performance is significantly worse than previous task-specific models and even for puzzles solvable by logic, the model is still not perfect potentially due to the complexity of applying the given chain of heuristics.
Another line of works investigates the reasoning capabilities of Masked Diffusion Models (MDMs) in language generation through the Sudoku task. (YGG+, 25) show the potential of MDMs to outperform autoregressive models in reasoning and planning with fewer parameters, highlighting their potential impact on downstream applications. (KSK+, 25) investigate the same models, exploring sampling strategies and token ordering effects, and propose an adaptive inference method that further improves performance. While (SDWP, 24) is more closely related to our work, our results are also comparable to those of (KSK+, 25), who outperformed (YGG+, 25). This demonstrates that our methodology surpasses all existing state-of-the-art approaches for solving combinatorial tasks such as Sudoku.
@@ 2 Methodology
In this work, we introduce a novel framework for reasoning with Transformers. Our method combines imitation learning with trial-and-error search, enabling models to go beyond learned heuristics. Specifically, when no further logical steps can be applied, the model is trained to make informed guesses, just as humans do. Crucially, it can backtrack when a guess eventually leads to a dead end, and attempt alternative continuations in a Depth-First Search (DFS) manner. By coupling simple learned rules with the flexibility to explore, our method navigates the solution space efficiently. The resulting Causal Transformer solves puzzles robustly and interpretably, showcasing emerging behaviors of progressive reasoning and self-correction.
@@
Our experiments showcase the capability of Transformers to learn this algorithmic behavior. Specifically, we observe that the models effectively learn to fill the values based on the Sudoku rules. When direct rule application is exhausted, the model guesses a candidate value and recovers correctly from incorrect guesses by backtracking when a dead end is identified, as in the DPLL algorithm (DP, 60; DLL, 62). As a result of these learned abilities, our method achieved near-perfect accuracy on Sudoku puzzles (99%), demonstrating state-of-the-art performance for a NN-based Sudoku solver (see Figure 1).
Our framework applies to any NP problem, whose solutions are hard to find but easy to verify. To demonstrate its generality, we also apply it to the canonical NP-complete problem of Satisfiability (Coo, 71), specifically the 1-in-3 SAT variant, where each clause contains three literals and exactly one must be true. Even in this task, our trained Transformer is able to correctly find solutions 99% of the time. Note that 1-in-3 SAT is a canonical NP-complete problem, and by the Cook-Levin theorem (Coo, 71), any problem in NP can be written as 1-in-3 SAT. However, some problems may benefit from problem-specific optimizations and therefore, we provide a library in Appendix F.2 that lets one generate transcripts for any problem in NP by providing the appropriate components.
@@
Whereas earlier studies struggled with accurately solving Sudoku, our approach can teach Transformers to solve combinatorial problems with near-perfect accuracy. Our thesis is that once LLMs can learn to solve different problems correctly, the next milestone is to also minimize the reasoning time. In fact, it is well-known that any combinatorial problem can be solved given sufficient time. As long as Transformers can reliably detect valid solutions to a problem, they can also solve it by randomly generating tokens and checking whether the produced transcript contains a solution. According to the “infinite-monkey” theorem (Mon, ), an infinitely large sequence of random tokens will almost-surely contain a valid solution to the problem instance.
@@
All our experiments follow a Causal Transformer (GPT-2 variant (RWC+, 19)) that is identical to that of (SDWP, 24) to allow for a direct comparison. The model comprises 8 layers, each with 8 self-attention heads, an embedding dimension of 576, and a feed-forward network with hidden dimension 3456 (6 $\times$ the embedding size). The model also employs causal masking within its attention layers to prevent access to future tokens. In total, this model contains 42M parameters.
We train and evaluate our model using randomly generated Sudoku. We deviate from previous works that used pre-existing datasets, such as the one provided by (Rad, 20), striving for a clean and interpretable distribution of Sudoku puzzles. In particular, our generator works exactly as follows:
@@ 3.1 Method overview
In (SDWP, 24), the authors train a GPT-2 variant to solve Sudoku puzzles using supervised imitation learning. Their training data consists of step-by-step solution transcripts that progressively apply seven human-crafted strategies (namely the Lone Single, Hidden Single, Naked Pair, Naked Triplet, Locked Candidate, XY Wing, and Unique Rectangle), ordered from simpler to more complex. For token encoding, each cell is represented using three tokens: ($r$, $c$, $v$), where $r$ and $c$ denote the cell’s row and column indices, and $v$ is the digit placed at that position.
@@
We adopt a novel tokenization strategy, encoding each move as a single 3-digit number ranging from 111 to 999, with the first digit representing the row r the second the column c and the third the value v. This differs from (SDWP, 24), that use 3 separate tokens for the row, column and value. Our action-level representation reduces the input sequence length by a factor of 3, leading to faster training iterations and more efficient solution generation during evaluation.
@@ 3.2 Benchmarks
We evaluate our model across four distinct datasets to benchmark its performance and ensure generalization: (i) Random: our own puzzle generator, used to produce both training data and a held-out 100K test set, (ii) Kaggle unfiltered: the Kaggle dataset (Rad, 20), a uniformly sampled collection of 3M puzzles, from which we randomly sample 100K for evaluation, (iii) Kaggle filtered: a 1.9M puzzle subset of the unfiltered Kaggle dataset, curated by (SDWP, 24) to include only puzzles solvable using seven Sudoku strategies, and (iv) RRN: the dataset introduced by (PPW, 18), consisting of 180K training and an 18K test set, stratified by clue count (ranging from 17 to 34 initial givens).
3.3 Results
We begin by evaluating how effectively the Transformer learns to apply basic Sudoku rules, and examine how our design choices, specifically the action-level token encoding and multi-target loss, affect the speed and efficiency of learning. In this setting, rule-logic accuracy is defined as the model’s ability to correctly fill all cells resolvable by rule-based inference. If even a single such cell is left unfilled or incorrectly filled, the entire board is considered incorrect. Figure 4 shows the evolution of rule-logic accuracy during training on our Random dataset, plotted against the number of tokens seen (in billions), which also correlates closely with training time. We compare three configurations: (i) the baseline setup from (SDWP, 24) using their triplet tokenization and single-target loss, (ii) our compact encoding with single-target loss, and (iii) our full approach with both compact encoding and multi-target loss. While all configurations eventually achieve high accuracy, our method substantially accelerates learning, with the multi-target variant reaching near-perfect performance much earlier.
@@ 3.3 Results
Causal Transformer (SDWP, 24) Kaggle filt. $\dagger 67.16$ 87.18% 94.80%
MDM (KSK+, 25) Kaggle filt. $\dagger 74.97$ 89.49% -$\ddagger$
@@
Regarding training, our model is trained using puzzles generated by our random generator. For the baselines, the Causal Transformer from (SDWP, 24) was trained on the Kaggle filtered dataset; the same holds for MDM (KSK+, 25). The RRN (PPW, 18) and Recurrent Transformer (YIL, 23) models were originally trained on the RRN dataset. For a fair comparison, we also trained these models on our random generator using their released training code.
The following key takeaways emerge from the comparative analysis in Table 2: first, our trial-and-error Transformer achieves significant improvements over the Causal Transformer baseline (SDWP, 24), with absolute accuracy gains of 31.7% (Kaggle unfiltered), 12.3% (Kaggle filtered), and 4.6% (RRN). Notably, while their model is restricted to puzzles solvable by seven human-crafted strategies (covering only 1.9M of Kaggle’s 3M puzzles), our method generalizes across the full distribution, achieving 98.9% accuracy on all 3M puzzles. Also, while domain-specific models like RRN and the Recurrent Transformer achieve strong performance on the RRN benchmark, our method outperforms them by 6-7% on the more diverse Random and Kaggle datasets, even when these competitors are enhanced by training on our random generator. For example, our model achieves 98.9% vs. their 92.4-92.6% on Random dataset. Finally we see that training on our random generator improves performance on diverse datasets (e.g., RRN’s +13.1% gain on Random test) while preserving RRN benchmark accuracy (97.9% vs. 98.9% when RRN-trained), demonstrating broad coverage without compromising specialization.

### v2

@@ Abstract
Despite their proficiency in various language tasks, Large Language Models (LLMs) struggle with combinatorial problems like Satisfiability, Traveling Salesman Problem, or even basic arithmetic. We address this gap through a novel trial & error approach for solving problems in the class NP, where candidate solutions are iteratively generated and efficiently validated using verifiers. We focus on the paradigmatic task of Sudoku and achieve state-of-the-art accuracy (99%) compared to prior neuro-symbolic approaches. Unlike prior work that used custom architectures, our method employs a vanilla decoder-only Transformer (GPT-2) without external tools or function calling. Our method integrates imitation learning of simple Sudoku rules with an explicit Depth-First Search (DFS) exploration strategy involving informed guessing and backtracking. Moving beyond imitation learning, we seek to minimize the number of guesses until reaching a solution. This is achieved using depth-1 guessing, showing empirically that almost all Sudoku can be solved using the puzzle’s rules with at most one guess. We provide a rigorous analysis of this setup formalizing its connection to a contextual variant of Min–Sum Set Cover, a well-studied problem in algorithms and stochastic optimization.
@@ 1 Introduction
Large Language Models (LLMs) based on the Transformer architecture have shown remarkable performance in a wide range of Machine Learning tasks, achieving state-of-the-art results on language understanding and generation benchmarks (Vaswani et al., 2017; Brown et al., 2020; Radford et al., 2019). Trained to predict the next token in large text corpora, these models acquire strong reasoning abilities, but whether they perform genuine logical reasoning or rely on pattern-matching remains under debate (Valmeekam et al., 2023; Dziri et al., 2023). Prompting strategies such as Chain-of-Thought and its many variants (Nye et al., 2021; Wei et al., 2022; Yao et al., 2023; Wang et al., 2023; Besta et al., 2024; Chen et al., 2023; Khot et al., 2022; Zhou et al., 2023; Madaan et al., 2023) have enabled LLMs to exhibit multi-step reasoning capabilities and solve even complex mathematical problems (Hendrycks et al., 2021). Yet, LLMs still struggle in more structured domains like arithmetic and combinatorial optimization. Controlled evaluations on synthetic tasks, including algorithmic puzzles, formal grammars, logic-based games, and arithmetic problems, have revealed consistent failure modes of current models and methods (Liu et al., 2022; Li et al., 2023; Nanda et al., 2023; Allen-Zhu and Li, 2023; Ye et al., 2024a, b; Lee et al., 2024). At the same time, the literature highlights that Transformer-based models have the potential to leverage
@@
internal mechanisms like attention to capture combinatorial structure without specialized heuristics or external tools. However, appropriately training them to display this behavior remains a major challenge and an active topic of research in robustness and generalization.
@@
Our primary testbed in this work is Sudoku, chosen for its well-defined rules, extensively studied difficulty levels, and widespread use as a reasoning benchmark. A standard Sudoku puzzle consists of a $9\times 9$ grid, subdivided into nine $3\times 3$ boxes. The grid is partially filled with digits from 1 to 9, and the goal is to complete it subject to the following rules: (i) each cell contains exactly one digit, (ii) each row contains every digit exactly once, (iii) each column contains every digit exactly once, and (iv) each box contains every digit exactly once. We focus on the standard Sudoku puzzles, while our approach can be directly applied to its generalized form with $n\times n$ grids which is NP-complete (Yato and Seta, 2003) and thus equivalent to any problem in the class NP.
Solving Sudoku puzzles without external tools is extremely challenging even for state-of-the-art industrial-scale LLMs. We evaluated several models, including OpenAI GPT (Achiam et al., 2023), Google Gemini (Team et al., 2023), and DeepSeek (Guo et al., 2025), on randomly generated Sudoku puzzles. As shown in Table 1, none of these models were able to produce correct solutions. We tested both general-purpose LLMs and reasoning-focused variants, since Sudoku demands logical reasoning. General-purpose models generated answers quickly, whereas reasoning-oriented models ran significantly longer as they attempted step-by-step strategies mirroring human solving patterns. However, once they made an incorrect inference, these models could not backtrack and revise their solutions, often ending with contradictory board states. Additional details are provided in Appendix B.
@@
Beyond these general-purpose models, several methods have been proposed in the literature to solve Sudoku puzzles via Deep Learning although often relying on specialized architectures. (Mladenov et al., 2011) investigates a Neural Network (NN)-based approach based on Hopfield Networks as an integer optimization tool, addressing its convergence limitations through a coprocessor NN designed to solve linear programming problems. (Palm et al., 2018) introduces the Recurrent Relational Network (RRN), a generalization of the Relational Network (Santoro et al., 2017), to evaluate multi-step relational reasoning. Their method involves multiple rounds of message passing between Sudoku puzzle cells to arrive at a solution. (Yang et al., 2023) explores solving Sudoku puzzles using a Recurrent Transformer model, integrating Sudoku’s constraints directly into both the model architecture and the training process.
@@
Closer to our work, (Shah et al., 2024) study how a causal language model (GPT-based) can reliably solve Sudoku puzzles. By reframing Sudoku as a next-token prediction task over transcripts of human-like reasoning based on known Sudoku strategies, their Causal Transformer learns to imitate and apply these heuristics incrementally, typically executing simpler strategies before tackling more complex ones. This approach showcases the promising capability of Autoregressive Transformers to successfully internalize and deploy multi-step reasoning procedures in a more interpretable way. However, their proposed method trains the Transformer on a structured sequence of logical steps, which do not allow the model to make mistakes. This inherently limits its problem-solving capability to strategies present in the training data whereas a large fraction of puzzles are not solvable by simple logic steps. As such, its performance is significantly worse than previous task-specific models and even for puzzles solvable by logic, the model is still not perfect potentially due to the complexity of applying the given chain of heuristics.
Another line of works investigates the reasoning capabilities of Masked Diffusion Models (MDMs) in language generation through the Sudoku task. (Ye et al., 2025) show the potential of MDMs to outperform autoregressive models in reasoning and planning with fewer parameters, highlighting their potential impact on downstream applications. (Kim et al., 2025) investigate the same models, exploring sampling strategies and token ordering effects, and propose an adaptive inference method that further improves performance. While (Shah et al., 2024) is more closely related to our work, our results are also comparable to those of (Kim et al., 2025), who outperformed (Ye et al., 2025). This demonstrates that our methodology surpasses all existing state-of-the-art approaches for solving combinatorial tasks such as Sudoku.
@@ 2 Methodology
In this work, we introduce a novel framework for reasoning with Transformers. Our method combines imitation learning with trial-and-error search, enabling models to go beyond learned heuristics. For combinatorial problems such as Sudoku, a verifier exists that can check whether a partial or complete assignment satisfies the problem’s rules; we therefore use this verifier to validate the correctness of each attempted instance. Specifically, when no further logical steps can be applied, the model is trained to make informed guesses, just as humans do. Crucially, it can backtrack when a guess eventually leads to a dead end, and attempt alternative continuations in a Depth-First Search (DFS) manner. By coupling simple learned rules with the flexibility to explore, while continuously relying on the verifier for correctness checks, our method navigates the solution space efficiently. The resulting Causal Transformer solves puzzles robustly and interpretably, showcasing emerging behaviors of progressive reasoning and self-correction.
@@
Our experiments showcase the capability of Transformers to learn this algorithmic behavior. Specifically, we observe that the models effectively learn to fill the values based on the Sudoku rules. When direct rule application is exhausted, the model guesses a candidate value and recovers correctly from incorrect guesses by backtracking when a dead end is identified, as in the DPLL algorithm (Davis and Putnam, 1960; Davis et al., 1962). As a result of these learned abilities, our method achieved near-perfect accuracy on Sudoku puzzles (99%), demonstrating state-of-the-art performance for a NN-based Sudoku solver (see Figure 1).
Our framework applies to any NP problem, whose solutions are hard to find but easy to verify. To demonstrate its generality, we also apply it to the canonical NP-complete problem of Satisfiability (Cook, 1971), specifically the 1-in-3 SAT variant, where each clause contains three literals and exactly one must be true. Even in this task, our trained Transformer is able to correctly find solutions 99% of the time. Note that 1-in-3 SAT is a canonical NP-complete problem, and by the Cook-Levin theorem (Cook, 1971), any problem in NP can be written as 1-in-3 SAT. However, some problems may benefit from problem-specific optimizations and therefore, we provide a library in Appendix F.2 that lets one generate transcripts for any problem in NP by providing the appropriate components.
@@
Whereas earlier studies struggled with accurately solving Sudoku, our approach can teach Transformers to solve combinatorial problems with near-perfect accuracy. Our thesis is that once LLMs can learn to solve different problems correctly, the next milestone is to also minimize the reasoning time. In fact, it is well-known that any combinatorial problem can be solved given sufficient time. As long as Transformers can reliably detect valid solutions to a problem, they can also solve it by randomly generating tokens and checking whether the produced transcript contains a solution. According to the “infinite-monkey” theorem (46), an infinitely large sequence of random tokens will almost-surely contain a valid solution to the problem instance.
@@
Despite its simplicity, this setup remains surprisingly effective, even outperforming the previous multi-level guessing method. This advantage arises because, in our beyond imitation learning approach, training uses only successful trajectories, i.e., those that contain the correct guess. While imitation learning provides richer supervision by including both correct and incorrect guess tokens, it also encourages the model through next-token prediction to continue generating incorrect tokens, rather than identifying the correct path. By excluding incorrect tokens from the training data, our approach goes beyond imitation learning and concentrates learning on high-value decisions that directly lead to a solution.
@@
All our experiments follow a Causal Transformer (GPT-2 variant (Radford et al., 2019)) that is identical to that of (Shah et al., 2024) to allow for a direct comparison. The model comprises 8 layers, each with 8 self-attention heads, an embedding dimension of 576, and a feed-forward network with hidden dimension 3456 (6 $\times$ the embedding size). The model also employs causal masking within its attention layers to prevent access to future tokens. In total, this model contains 42M parameters.
We train and evaluate our model using randomly generated Sudoku. We deviate from previous works that used pre-existing datasets, such as the one provided by (Radcliffe, 2020), striving for a clean and interpretable distribution of Sudoku puzzles. In particular, our generator works exactly as follows:
@@ 3.1 Method overview
In (Shah et al., 2024), the authors train a GPT-2 variant to solve Sudoku puzzles using supervised imitation learning. Their training data consists of step-by-step solution transcripts that progressively apply seven human-crafted strategies (namely the Lone Single, Hidden Single, Naked Pair, Naked Triplet, Locked Candidate, XY Wing, and Unique Rectangle), ordered from simpler to more complex. For token encoding, each cell is represented using three tokens: ($r$, $c$, $v$), where $r$ and $c$ denote the cell’s row and column indices, and $v$ is the digit placed at that position.
@@
We adopt a novel tokenization strategy, encoding each move as a single 3-digit number ranging from 111 to 999, with the first digit representing the row r the second the column c and the third the value v. This differs from (Shah et al., 2024), that use 3 separate tokens for the row, column and value. Our action-level representation reduces the input sequence length by a factor of 3, leading to faster training iterations and more efficient solution generation during evaluation.
@@ 3.2 Benchmarks
We evaluate our model across four distinct datasets to benchmark its performance and ensure generalization: (i) Random: our own puzzle generator, used to produce both training data and a held-out 100K test set, (ii) Kaggle unfiltered: the Kaggle dataset (Radcliffe, 2020), a uniformly sampled collection of 3M puzzles, from which we randomly sample 100K for evaluation, (iii) Kaggle filtered: a 1.9M puzzle subset of the unfiltered Kaggle dataset, curated by (Shah et al., 2024) to include only puzzles solvable using seven Sudoku strategies, and (iv) RRN: the dataset introduced by (Palm et al., 2018), consisting of 180K training and an 18K test set, stratified by clue count (ranging from 17 to 34 initial givens).
@@
3.3 Results
We begin by evaluating how effectively the Transformer learns to apply basic Sudoku rules, and examine how our design choices, specifically the action-level token encoding and multi-target loss, affect the speed and efficiency of learning. In this setting, rule-logic accuracy is defined as the model’s ability to correctly fill all cells resolvable by rule-based inference. If even a single such cell is left unfilled or incorrectly filled, the entire board is considered incorrect. Figure 4 shows the evolution of rule-logic accuracy during training on our Random dataset, plotted against the number of tokens seen (in billions), which also correlates closely with training time. We compare three configurations: (i) the baseline setup from (Shah et al., 2024) using their triplet tokenization and single-target loss, (ii) our compact encoding with single-target loss, and (iii) our full approach with both compact encoding and multi-target loss. While all configurations eventually achieve high accuracy, our method substantially accelerates learning, with the multi-target variant reaching near-perfect performance much earlier.
@@ 3.3 Results
Causal Transformer (Shah et al., 2024) Kaggle filt. $\dagger 67.16$ 87.18% 94.80%
MDM (Kim et al., 2025) Kaggle filt. $\dagger 74.97$ 89.49% -$\ddagger$
@@
Regarding training, our model is trained using puzzles generated by our random generator. For the baselines, the Causal Transformer from (Shah et al., 2024) was trained on the Kaggle filtered dataset; the same holds for MDM (Kim et al., 2025). The RRN (Palm et al., 2018) and Recurrent Transformer (Yang et al., 2023) models were originally trained on the RRN dataset. For a fair comparison, we also trained these models on our random generator using their released training code.
The following key takeaways emerge from the comparative analysis in Table 2: first, our trial-and-error Transformer achieves significant improvements over the Causal Transformer baseline (Shah et al., 2024), with absolute accuracy gains of 31.7% (Kaggle unfiltered), 12.3% (Kaggle filtered), and 4.6% (RRN). Notably, while their model is restricted to puzzles solvable by seven human-crafted strategies (covering only 1.9M of Kaggle’s 3M puzzles), our method generalizes across the full distribution, achieving 98.9% accuracy on all 3M puzzles. Also, while domain-specific models like RRN and the Recurrent Transformer achieve strong performance on the RRN benchmark, our method outperforms them by 6-7% on the more diverse Random and Kaggle datasets, even when these competitors are enhanced by training on our random generator. For example, our model achieves 98.9% vs. their 92.4-92.6% on Random dataset. Finally we see that training on our random generator improves performance on diverse datasets (e.g., RRN’s +13.1% gain on Random test) while preserving RRN benchmark accuracy (97.9% vs. 98.9% when RRN-trained), demonstrating broad coverage without compromising specialization.